# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v1

@@ 3.2.3 Profile Module
Similarly to the approach presented by the paper [4], our framework defines the function of the profile module as to establish the role of the LLM and adopts more diverse methods. This template explicitly separates the roles of both the core-agent and the LLM. Indeed, the profile module facilitates the dynamic adaptation of various profiles tailored to specific use cases and strategies employed by the planning module.
The Profile module features four methods for defining profiles: the Handcrafted In-Context Learning Method, the LLM-generation method, the Dataset Alignment Method, and the newly introduced Fine-tuned Pluggable Modules method.
@@
Finally, we introduced the Fine-tuned Pluggable Modules Method, a pioneering solution to set LLM’s profile leveraging several state-of-the-art techniques, aiming to provide an efficient customization and adaptation of the LLM’s profile. This approach defines the LLM profile by injecting a pluggable module fine-tuned to influence the language model behavior. Indeed, that module is a set of additional tunable parameters that must be previously trained using Parameter-Efficient Fine-Tuning (PEFT) techniques, such as Sequential adapter (AdapterS), Prompt-tuning, or LoRA [35].
This process ensures precise and effective customization to achieve the desired profile by eliminating the need for in-context learning, which traditionally involves injecting extensive prompt information into each query. By omitting this step, the required context size is significantly reduced, as the adapters directly encode the desired behaviors and profiles into the model parameters. This reduction in context size and memory footprint allows the LLM to operate more efficiently, using less memory during inference without compromising performance.
@@ 3.2.4 Action Module
While we have enriched the approaches applied in this module, we define it within the LLM-Agent-UMF similarly to the methodology described in [4]. This module serves as the convergence point where contributions from all other modules and their latent inputs are visualized as outcomes. It translates the agent’s decisions into executable actions, conceptualized from four perspectives: Action Goal, Action Production, Action Space, and Action Impact.
The first perspective, the Action Goal, represents what the core-agent intends to accomplish by performing actions based on various objectives such as Task Completion, Communication, or Environment Exploration. The Action module maintains this Action Goal and guides the core-agent in selecting appropriate actions accordingly. The second aspect, Action Production, focuses on how the actions are produced and the catalysts behind them. In fact, the core-agent can produce actions through three primary methods: Action via Memory Recollection, Action via Plan Following, and Action via API Call Request, newly introduced in our framework.
In the case of Action via Memory Recollection, the core-agent generates actions by extracting relevant information from its memory based on the current task. For instance, in GITM, if the agent needs to achieve a sub-goal, it checks its memory for similar past experiences and uses successful actions from those experiences to handle the current task [31].
Similarly, through the Action via Plan Following method, the core-agent can execute actions by adhering to pre-generated plans. In GITM and DEPS (Describe, Explain, Plan, and Select) [36], the agent breaks down the task into subtasks and sub-procedures, acting sequentially to achieve each procedure and complete the overall task. In case of plan failure, both agents have replanning capabilities, so the action module executes the new plan. These two methods, Action via Memory Recollection and Action via Plan Following, align with the framework described in [4].
By exploring other state-of-the-art techniques like TALM [5] and ToolFormer [37] that exemplify LLMs’ capacity to improve performance across diverse tasks through incorporating external tools, we introduce a new method named Action via API Call Request. This novel approach empowers the core-agent to execute actions in response to API call requests initiated by the LLM, facilitating smooth integration and effective utilization of external resources.
@@
The third perspective, Action Space, defines the set of possible actions that can be performed by the core-agent, remaining consistent with the original framework [4]. These actions can leverage internal knowledge, where the core-agent acts on internal information, and/or utilizes tools when comprehensive expert knowledge is required. These tools can encompass APIs such as HuggingGPT, which leverages AI models via the HuggingFace API to accomplish complex user tasks [38]. Alternatively, the core-agent can rely not just on its memory but extend its operational scope and knowledge by communicating with other read-only data sources like databases and knowledge repositories. It is useful to note here that data repositories can have the same formats discussed in the memory module (section 3.2.2). For instance, LLamaIndex [39] stores data as vector embeddings at the indexing stage to leverage semantic search, where the similarity between embeddings is used to rank documents by their relevance to a query. In contrast, ReAct [40] uses a textual data repository, like Wikipedia, to mitigate error propagation in chain-of-thought reasoning.
@@
Finally, Action impact refers to the consequences resulting from an action. Numerous impacts can be cited, such as changing the environment by moving to different locations, gathering resources, or constructing buildings [41]. Actions can also alter the internal state of the core-agent by updating its memory or acquiring new knowledge which may affect the planning module. Additionally, impacts can trigger new actions, creating a chain of actions during task completion.
@@ 3.2.5 Security Module
As LLMs continue to advance and being implemented in widespread application, addressing the potential risks and unintended consequences associated with their use becomes imperative. These concerns are mainly around unauthorized and unethical use, data biases and privacy [3] . This has led to the introduction of guardrails recently in LLMs field as algorithms to identify and prevent the misuse of LLMs [12]. To enhance our framework, we propose a fifth module: the Security Module. This module aims to provide a more capable and responsible core-agent. Its role is monitoring the action module specifically in production environments to ensure the safety and responsible use of LLMs.
@@
While conducting a thorough research around the Security Module, our approach will involve exploring multiple facets: Identifying and securing critical assets and data within the core-agent modeling framework, implementing strategies and mechanisms to protect these assets from potential threats, ensuring the core-agent’s ability to effectively respond to and mitigate any security incidents, and maintaining the privacy of user data while adhering to relevant data protection regulations.
@@
    The recent survey [47] has demonstrated that despite the implementation of prompt safeguarding techniques, the overall resilience of Large Language Models (LLMs) against advanced attacks, known as jailbreaks, may not experience significant improvement. These jailbreaks are designed to exploit biases or vulnerabilities within language models by manipulating their responses. Notable examples include white-box attacks AutoDAN-Zhu [48], which generate stealthy prompts to avoid triggering the model protective mechanisms. Additionally, black-box attacks leverage manually crafted prompts to deceive the LLM [47]. The effectiveness of these jailbreak techniques is further illustrated in [12], where researchers successfully achieved a jailbreak attack on ChatGPT 3.5 by framing potentially harmful query, “how to hotwire a car” as a hypothetical scenario. The existence of these jailbreak methods highlights the urgent necessity for continuous and rigorous monitoring of LLM outputs to detect and mitigate potential breaches.
    Indeed, guarding the agent outputs involves ensuring the security and integrity of the text generated by LLMs. This includes detecting and redacting harmful content while maintaining coherence and relevance. Two notable examples of such techniques are Guardrails AI [47] and LLMSafeGuard [49], discussed further in section 4.
@@ 3.3 Active/Passive Core-Agent Classification
As discussed in the preceding sections, the core-agent represents a distinct entity within the LLM-Agent-UMF. While the LLM excels in cognitive tasks such as understanding, reasoning, and generating responses, it lacks the capability to directly interact with the environment or external tools. This is where the core-agent plays a crucial role. It bridges the gap between the LLM’s cognitive abilities and the need to engage with external sources, enabling seamless integration with various tools and systems. The core-agent is thus characterized by its action capabilities and its ability to respond to user requests through interaction with these diverse tools. In fact, ToolLLM [54] is a general tool-use framework that enhances LLMs capabilities enabling agents to use external tools and APIs. It uses a neural API retriever to recommend appropriate APIs for each instruction. Then they employ a depth-first search-based decision tree algorithm to evaluate multiple reasoning traces and expand the search space. Consequently, it enhances the planning ability of the retriever and empowers the finetuned LLM, ToolLlaMA, to generate adequate instructions. The retriever here, in association with the search-based decision tree algorithm, satisfies our definition of a core-agent. In this case where the LLM-based agent conducts cognitive tasks, memory and planning modules are essential in the core-agent to ensure reasoning capabilities because they enable the agent to retain and recall past experiences, plan and synchronize actions, reason and make decisions [26] [21].
@@
The inspection of the state-of-the-art led to the conclusion that the action module is always indispensable in a core-agent as it is responsible for producing the executive steps to achieve their goals. However, the architectural disparities in core-agents and the absence of some modules in some proposed agent systems highlight the need to introduce a new taxonomy classifying core-agents into two distinct categories: Active core-agents and passive core-agents. The following sections analyze and pinpoint the main differences and similarities between active and passive core-agents in their structural alignment with our framework.
@@ 3.3.1 Active Core-Agents
3.3.2 Passive Core-Agents
Table 1: Active and passive core-agent internal structure
@@ 3.3.2 Passive Core-Agents
Handcrafted in-context learning method X [*]
Fine-tuned pluggable modules method X [*]
LLM-generation method X
Profile Dataset alignment method X
Embedded memory
Memory location Memory extension
Short-term
Memory structure Long-term
Natural language
SQL database
Embeddings
Memory Memory format Structured list X
@@
Planning Feedback sources Tools feedback X
@@
Action via Memory Recollection X
Action via Plan Following X X
Action production Action via API Call Request X X
Internal knowledge X X
Action space Tools (APIs, Data repositories, External systems, etc) X X
@@
Rule-based guardrail
@@
Reinforce Single Responsibility Principle (SRP). X
@@
Advantages Protects against adversarial attacks. X
@@
Synchronization is needed in multi core-agents’ setups. X
@@
Preclude human-initiated communication. X
Disadvantages Lacks memory, limiting visibility into the agent’s overall status. X
@@ 3.4.1 Uniform multi-core agent
  • Uniform multi-passive core-agent architecture leverages passive core-agents’ capabilities of handling low-level operations and executing specific tasks. The configuration shown in Figure 5 is an example where the LLM communicates with multiple passive core-agents and harnesses strategically their singular strengths to retrieve diverse information or perform specialized functions to generate a comprehensive final output. In fact, the language model assumes leadership and complete control over the ensemble of passive core-agents.
    In essence, uniform multi-passive core-agent systems are distinguished by the ease of integration of new passive core-agents, thereby extending their functionality without the need for complex synchronization. Consequently, the only modification resulting from the introduction of new passive core-agents involves adapting the LLM profile, either statically at setup/configuration time or dynamically via an active core-agent, as will be discussed in the context of the hybrid setup.
@@
  • Uniform active core-agents architecture deals with the interaction of a group of active core-agents in one system as illustrated in Figure 6. As opposite to passive core-agents that operate solely with action and security modules, active core-agents possess all five modules (Planning, Memory, Profile, Action and Security) enabling them to manage complex cognitive tasks. This architecture may be seen as a better alternative to uniform multi-passive core-agents’ design due to its wider range of capabilities and functionalities. However, due to the authoritative nature of the active entities, the multi-active core-agent design is more complex than the one exclusively based on passive core-agents. The inclusion of multiple active elements in one system introduces challenges similar to those in multi-agent systems. For instance, effective communication among active core-agents is paramount; given their dynamic nature, timely and accurate exchange of information _such as inter-sibling core-agent feedback and status updates_ is crucial to ensure cohesive operation of the agent. As the number of active core-agents grows, managing intra-communication becomes increasingly complex resulting in frequently emerging synchronization issues. This is why multi-active core-agent systems potentially necessitate consensus algorithms, such as Raft [56] [57], to elect a leader.
    Therefore, we deduce that the independent implementation of either of these architectures is limited and may be problematic. While the multi-passive core-agents architecture is efficient in granular task execution and low-level operations, it lacks a component for handling high-level tasks such as decision-making, task planning, and resource allocation which are intrinsic to multi-active core-agent systems. However, the latter introduces synchronization issues and increases system complexity. This dilemma compels us to introduce the hybrid approach.
@@ 3.4.2 Hybrid multi-core agent
To leverage the strength of both passive and active core-agent architectures, we propose an optimal system depicted in Figure 7. It integrates one active entity as the manager with multiple passive entities functioning as workers within a unified system. One managerial aspect of the active core-agent is its ability to configure dynamically the profile of LLMs, enabling them to effectively utilize passive core-agents for handling specific tasks. This configuration leverages the parallel execution capabilities of numerous passive core-agents under the guidance and leadership of an active core-agent, empowering the system to handle wider range of tasks, while preserving a comfortable level of flexibility, extendibility and scalability.
This hybrid design realizes the full potential of the multi-core architecture: by uniting the strengths of the uniform passive core-agent architecture with the capabilities of the active core-agent, the system can dynamically allocate resources and adjust its configuration based on the specific requirements of the task at hand. Our proposed architecture of one-active-many-passive strikes a balance between the intricate nature of multi-active core-agent architectures and the practicality offered by passive core-agents.
Figure 7: One-active-many-passive core-agent hybrid architecture
@@
Figure 8: Many-active-many-passive core-agent hybrid architecture

### v2

@@ 3.2.3 Profile Module
Similarly to the approach presented by the paper [5], our framework defines the function of the profile module as to establish the role of the LLM and adopts more diverse methods. This template explicitly separates the roles of both the core-agent and the LLM. Indeed, the profile module facilitates the dynamic adaptation of various profiles tailored to specific use cases and strategies employed by the planning module. The Profile module features four methods for defining profiles: the Handcrafted In-Context Learning Method, the LLM-generation method, the Dataset Alignment Method, and the newly introduced Fine-tuned Pluggable Modules method (Figure 6).
Figure 6: Profile module techniques
@@
Finally, we introduced the Fine-tuned Pluggable Modules Method, a pioneering solution to set LLM’s profile leveraging several state-of-the-art techniques, aiming to provide an efficient customization and adaptation of the LLM’s profile. This approach defines the LLM profile by injecting a pluggable module fine-tuned to influence the language model behavior [36]. Indeed, that module is a set of additional tunable parameters that must be previously trained using Parameter-Efficient Fine-Tuning (PEFT) techniques, such as Sequential adapter (AdapterS), Prompt-tuning, or LoRA [37].
This process ensures precise and effective customization to achieve the desired profile by eliminating the need for in-context learning, which traditionally involves injecting extensive prompt information into each query. By omitting this step, the required context size is significantly reduced, as the adapters directly encode the desired behaviors and profiles into the model parameters. This reduction in context size and memory footprint allows the LLM to operate more efficiently, using less memory during inference without compromising performance and flexibility.
@@ 3.2.4 Action Module
While we have enriched the approaches applied in this module, we define it within the LLM-Agent-UMF almost similarly to the methodology described in [5]. The action module is responsible for converting high-level instructions from the planning module into low-level actions, leveraging tools available in its environment. In our framework, the action module interacts with the security module to ensure that any executed action aligns with predefined criteria and prevent information leakage as will be explained further in Section 3.2.5. In essence, our analysis considers four fundamental aspects: Goal, Trigger, Action Space, and Impact (Figure 7).
Figure 7: Action module functional perspectives
The first perspective, the Action Goal, represents what the core-agent intends to accomplish by performing actions based on various objectives such as Task Completion, Communication, or Environment Exploration which aligns with the framework proposed in [5]. The second aspect, Action Trigger, focuses on how the actions are produced and the catalysts behind them. This perspective is derived from the Action Production perspective originally proposed in [5], but we emphasize more on the causality that drives action production. In this regard, the "Action via Memory Recollection" approach mentioned in [5] is considered to be better handled by the planning module as it focuses on leveraging past experiences for making appropriate decisions. In fact, the core-agent can produce actions as a result of two primary triggers: Plan Following and API Call Request, newly introduced in our framework. In the case of Plan Following trigger, the core-agent can execute actions by adhering to pre-generated plans elaborated by the planning module. As a matter of fact, in GITM [32], the agent breaks down the task into subtasks and procedures and executes the appropriate actions in the Minecraft world via the LLM interface. By exploring other state-of-the-art techniques like TALM [6] and ToolFormer [38] that exemplify LLMs’ capacity to improve performance across diverse tasks through incorporating external tools, we introduce a new trigger named API Call Request. This approach enables the core-agent to execute actions in response to API call requests initiated by the LLM, facilitating smooth integration and effective utilization of external resources.
@@
The third perspective, Action Space, defines the set of possible actions that can be performed by the core-agent in response to its objectives and environmental factors. This concept is distinct from operations involving internal state changes and memory management, which are handled separately by the Memory Module. In contrast to [5], our framework emphasizes this separation to avoid potential overlap between action and memory modules functions, thus ensuring a clear division of responsibilities for efficient execution. The action module can leverage tools ranging from basic software components like calculators to more advanced systems accessed through API calls, as exemplified in the HuggingGPT work that leverages the HuggingFace API to accomplish complex user tasks [39]. Alternatively, the core-agent may expand its operational scope and knowledge by communicating with read-only data sources such as knowledge repositories. This process is often referred to as Retrieval-Augmented Generation (RAG) [40]. This is particularly crucial for AI systems deployed in critical industries like healthcare where explainability in decision-making processes becomes paramount [41, 42]. By interacting with these external resources, transparency, provenance and traceability are ensured to guarantee the reliability of the agent’s outputs [43]. It is useful to note here that data repositories can have the same formats discussed in the memory module (Section 3.2.2). For instance, LLamaIndex [44] stores data as vector embeddings at the indexing stage to leverage semantic search, where the similarity between embeddings is used to rank documents by their relevance to a query. In contrast, ReAct [45] uses a textual data repository, like Wikipedia, to mitigate error propagation in chain-of-thought reasoning.
@@
Finally, Action Impact refers to the consequences resulting from an action. Numerous impacts can be cited, such as changing the environment by moving to different locations, gathering resources, or synthesizing new chemical compounds [12]. Actions can also result in the alteration of the internal state of the agent, especially when acquiring new knowledge and collaborating with the memory module. Additionally, actions can trigger new ones, creating a chain of actions during task execution.
@@ 3.2.5 Security Module
As LLMs continue to advance and become increasingly prevalent in various applications, it becomes crucial to address potential risks, ethical considerations and unintended consequences associated with their deployment [46]. These concerns revolve around issues such as unauthorized or unethical use, data biases, privacy breaches, and the spread of misinformation [4]. This has led to the introduction of guardrails recently in LLMs field as algorithms to identify and prevent the misuse of LLMs [13]. To bolster our framework for trustworthy AI systems, we propose a fifth module: the Security Module. This module aims to provide a more capable and responsible core-agent. Its role is monitoring the action module specifically in production environments to ensure the safety and responsible use of LLMs.
@@
Figure 8: Security module functional perspectives
While conducting a thorough research around the Security Module, our approach will involve exploring multiple facets: Identifying and securing critical assets and data within the core-agent modeling framework, implementing strategies and mechanisms to protect these assets from potential threats, ensuring the core-agent’s ability to effectively respond to and mitigate any security incidents, and maintaining the privacy of user data while adhering to relevant data protection regulations (Figure 8).
@@
    The recent survey [52] has demonstrated that despite the implementation of prompt safeguarding techniques, the overall resilience of Large Language Models (LLMs) against advanced attacks, known as jailbreaks, may not experience significant improvement. These jailbreaks are designed to exploit biases or vulnerabilities within language models by manipulating their responses. Notable examples include white-box attacks AutoDAN-Zhu [53], which generate stealthy prompts to avoid triggering the model protective mechanisms. Additionally, black-box attacks leverage manually crafted prompts to deceive the LLM [52]. The effectiveness of these jailbreak techniques is further illustrated in [13], where researchers successfully achieved a jailbreak attack on ChatGPT 3 by framing potentially harmful query, "how to hotwire a car" as a hypothetical scenario. The existence of these jailbreak methods highlights the urgent necessity for continuous and rigorous monitoring of LLM outputs to detect and mitigate potential breaches. It is useful to note that the aforementioned jailbreak attack was addressed and resolved in later versions of ChatGPT, 3.5 and 4, as confirmed by [13].
    Guarding the agent outputs is considered a supplementary measure alongside prompt safeguarding, with the objective of guaranteeing the safety, integrity and authenticity of the text generated by LLMs. This includes detecting and redacting harmful content while maintaining coherence and relevance. Two notable examples of such techniques are Guardrails AI [52] and LLMSafeGuard [54], discussed further in Section 4.
@@ 3.3 Active/Passive Core-Agent Classification
As discussed in the preceding sections, the core-agent represents a distinct entity within the LLM-Agent-UMF. While the LLM excels in cognitive tasks such as understanding, reasoning, and generating responses, it lacks the capability to directly interact with the environment or external tools. This is where the core-agent plays a crucial role. It bridges the gap between the LLM’s cognitive abilities and the need to engage with external sources, enabling seamless integration with various tools and systems. The core-agent is thus characterized by its action capabilities and its ability to respond to user requests through interaction with these diverse tools. In fact, ToolLLM [59] is a general tool-use framework that enhances LLMs capabilities enabling agents to use external tools and APIs. It uses a neural API retriever to recommend appropriate APIs for each instruction. Then they employ a depth-first search-based decision tree algorithm to evaluate multiple reasoning traces and expand the search space. Consequently, it enhances the planning ability of the retriever and empowers the finetuned LLM, ToolLlaMA, to generate adequate instructions. The retriever here, in association with the search-based decision tree algorithm, satisfies our definition of a core-agent. In this case where the LLM-based agent conducts cognitive tasks, memory and planning modules are essential in the core-agent to ensure reasoning capabilities because they enable the agent to retain and recall past experiences, plan and synchronize actions, reason and make decisions [27, 22].
@@
The inspection of the state-of-the-art led to the conclusion that the action module is always indispensable in a core-agent as it is responsible for producing the executive steps to achieve their goals. However, the architectural disparities in core-agents and the absence of some modules in some proposed agent systems highlight the need to introduce a new taxonomy classifying core-agents into two distinct categories: Active core-agents (Figure 3) and passive core-agents (Figure 9). The following sections analyze and pinpoint the main differences and similarities between active and passive core-agents in their structural alignment with our framework.
@@ 3.3.1 Active Core-Agents
Table 1: Active and passive core-agents internal structure
@@
Tools feedback
Planning Feedback sources Sibling core-agent feedback X
Embedded memory
Memory location Memory extension
Short-term
Memory scope Long-term
Natural language
SQL database
Embeddings
Memory Memory format Structured list X
Handcrafted in-context learning method X [*]
Fine-tuned pluggable modules method X [*]
LLM-generation method X
Profile Dataset alignment method X
@@
Plan Following X
Action trigger API Call Request X X
Tools (APIs, External systems, etc) X X
Action space Data repositories X X
@@
Rule-based guardrail X X
@@ 3.3.2 Passive Core-Agents
Advantages Reinforce Single Responsibility Principle (SRP). X
@@
Protects against adversarial attacks. X
@@
Disadvantages Synchronization is needed in multi core-agents’ setups. X
@@
Preclude human-initiated communication X
Lacks memory, limiting visibility into the agent’s overall status. X
@@ 3.4.1 Uniform multi-core agent
  • Uniform multi-passive core-agent architecture leverages passive core-agents’ capabilities of handling low-level operations and executing specific tasks. The configuration shown in Figure 10 is an example where the LLM communicates with multiple passive core-agents and harnesses strategically their singular strengths to retrieve diverse information or perform specialized functions to generate a comprehensive final output. In fact, the language model assumes leadership and complete control over the ensemble of passive core-agents. In essence, uniform multi-passive core-agent systems are distinguished by the ease of integration of new passive core-agents, thereby extending their functionality without the need for complex synchronization. Consequently, the only modification resulting from the introduction of new passive core-agents involves adapting the LLM profile, either statically at setup/configuration time or dynamically via an active core-agent, as will be discussed in the context of the hybrid setup.
  • Uniform active core-agents architecture deals with the interaction of a group of active core-agents in one system as illustrated in Figure 11. As opposite to passive core-agents that operate solely with action and security modules, active core-agents possess all five modules (Planning, Memory, Profile, Action and Security) enabling them to manage complex cognitive tasks. This architecture may be seen as a better alternative to uniform multi-passive core-agents’ design due to its wider range of capabilities and functionalities. However, due to the authoritative nature of the active entities, the multi-active core-agent design is more complex than the one exclusively based on passive core-agents. The inclusion of multiple active elements in one system introduces challenges similar to those in multi-agent systems. For instance, effective communication among active core-agents is paramount; given their dynamic nature, timely and accurate exchange of information _such as inter-sibling core-agent feedback and status updates_ is crucial to ensure cohesive operation of the agent. As the number of active core-agents grows, managing intra-communication becomes increasingly complex resulting in frequently emerging synchronization issues. This is why multi-active core-agent systems potentially necessitate consensus algorithms, such as Raft [61, 62], to elect a leader.
@@
    Therefore, we deduce that the independent implementation of either of these architectures is limited and may be problematic. While the multi-passive core-agent architecture is efficient in granular task execution and low-level operations, it lacks a component for handling high-level tasks such as decision-making, task planning, and resource allocation which are intrinsic to multi-active core-agent systems. However, the latter introduces synchronization issues and increases system complexity. This dilemma compels us to introduce the hybrid approach.
@@ 3.4.2 Hybrid multi-core agent
To leverage the strength of both passive and active core-agents architectures, we propose an optimal system depicted in Figure 12. It integrates one active entity as the manager with multiple passive entities functioning as workers within a unified system. One managerial aspect of the active core-agent is its ability to configure dynamically the profile of LLMs, enabling them to effectively utilize passive core-agents for handling specific tasks. This configuration leverages the parallel execution capabilities of numerous passive core-agents under the guidance and leadership of an active core-agent, empowering the system to handle wider range of tasks, while preserving a comfortable level of flexibility, extendibility and scalability.
Figure 12: One-active-many-passive core-agent architecture
This hybrid design realizes the full potential of the multi-core architecture: by uniting the strengths of the uniform passive core-agents architecture with the capabilities of the active core-agent, the system can dynamically allocate resources and adjust its configuration based on the specific requirements of the task at hand. Our proposed architecture of one-active-many-passive strikes a balance between the intricate nature of multi-active core-agent architectures and the practicality offered by passive core-agents.
@@
Figure 13: Many-active-many-passive core-agent architecture